OpenStudy (anonymous):

solve 2sin^2X=4sinX+6

OpenStudy (anonymous):

2sin^2X=4sinX+6 \[2 \sin 2x = 4 \sin x + 6\] it looks like this, isn't it?

OpenStudy (anonymous):

It is 2sin to the power of 2 x = 4sinx+6

OpenStudy (anonymous):

ahhh

OpenStudy (anonymous):

\[2 (\sin x )^2 = 4 \sin x + 6\]

OpenStudy (anonymous):

like this ?

OpenStudy (anonymous):

we assume sin x to be Y so we have \[2 Y^2 = 4Y + 6\]

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

\[2Y^2 = 4Y + 6\] \[2Y^2 - 4Y - 6 = 0\] \[y^2 - 2Y - 3 = 0\] \[(Y+1) (Y-3) = 0\]

OpenStudy (anonymous):

\[2\sin ^{2}x=4sinx+6\]

OpenStudy (anonymous):

get it ?

OpenStudy (anonymous):

it is like this \[2\sin ^{2}x=4sinx+6\]

OpenStudy (anonymous):

yess then we assume sin x to be Y like what i said before

OpenStudy (anonymous):

\[Y = -1\] because of Y equal to sin x, then \[\sin x = -1\] x = \[x = \sin^{-1} (-1) = -\frac{ \pi }{ 2 } \]

OpenStudy (anonymous):

let us prove it

OpenStudy (anonymous):

We have \[X = -\frac{ \pi }{ 2 }\] substitute this X to \[2 \sin^{2}x = 4 \sin x + 6\] then we have \[2 \sin^{2}(-\frac{ \pi }{ 2 }) = 4 \sin (-\frac{ \pi }{ 2 }) + 6\] 2 (-1)^2 = 4 (-1) + 6 2 = -4 + 6 2 = 2 Of course we have proved it! :)